# Always Keep Your Promises: DynamicLRP, A Model-Agnostic Solution To Layer-Wise Relevance Propagation

### v1

@@ Appendix E Input Aggregation
Remark. This property ensures that the total number of nodes traversed by all Promises is bounded by $\delta=\sum_{v_{p}\in V_{P}}d(v_{p})\leq|V|$, as stated in Theorem 1 (4.4). Each node contributes to at most one Promise’s internal chain, preventing double-counting in complexity analysis.
Appendix F Attribution Faithfulness Evaluation Figures
Figure 5: Comparing methods for VGG attributions.
Figure 6: Comparing faithfulness of methods using ABPC from pretrained ViT-b-16 performance on 1000 examples from the ImageNet CIFAR-10 test set. Occlusion is applied by replacing 16x16 patches with the corresponding regions from a Gaussian-blurred version of the image (kernel size = 51, $\sigma$ = 20). Red = MoRF curve, Blue = LeRF curve.

### v3

@@ Appendix E Input Aggregation
Remark. This property ensures that the total number of nodes traversed by all Promises is bounded by $\delta=\sum_{v_{p}\in V_{P}}d(v_{p})\leq|V|$, as stated in Theorem 1 (1). Each node contributes to at most one Promise’s internal chain, preventing double-counting in complexity analysis.
Appendix F List of Covered Operations
Below is the list of currently supported operations for DynamicLRP.
Category Operations
Arithmetic Add, Sub, Mul, Div, Neg, Sqrt, Rsqrt, Pow, Exp, Log
Linear Algebra Mm, Bmm, Convolution
Aggregation Sum, Mean, Cat, Stack, Unbind, Split, Max, Gather
Indexing and Selection Slice, Index, Select
Pooling MaxPool2D, MaxPool3D, AdaptiveAvgPool2D
Normalization NativeLayerNorm, NativeBatchNorm
Attention SDPA
Activations GELU, SiLU, ReLU, Softmax
Other MaskedFill, Clone, ToCopy, NativeDropout, Embedding
Appendix G Extra Results
Table A.1: ABPC and attribution efficiency metrics for VGG experiments. 3800 examples from ImageNette (320px) validation set.
Saliency 0.73 2.58 1.85 156.83 0.89GB
SmoothGrad 1.50 2.87 1.37 8.90 0.90GB
Random -0.0022 2.40 2.40 - -
Zennit LRP 1.69 2.72 1.02 29.41 1.4GB
Ours 1.77 2.65 0.88 21.79 2.3GB
Table A.2: ABPC and attribution efficiency metrics for ViT-b-16 experiments. 1000 examples from CIFAR-10 test set.
IG 0.10 0.95 0.85 N/A* >12GB
InputXGrad -0.0030 0.90 0.90 59.89 2.1GB
GradSHAP 0.074 0.94 0.87 3.32 8.6GB
Random -0.023 0.88 0.90 - -
AttnLRP 1.46 1.78 0.33 6.02 0.96GB
Ours 1.46 1.78 0.32 9.74 1.6GB
Figure A.1: Comparing methods for VGG attributions.
Table A.3: DynamicLRP Attribution faithfulness measured by top attributed token accuracy against model predictions and labels, and IoU with predictions with the entire SQuADv2 validation set (5928 examples). EM = Exact Match. Higher is better for all but the last two columns. Columns marked with (P) are w.r.t. model predictions, (L) are w.r.t. labels. Top-2 accuracy and IoU values in parentheses signify strict matching to only the predicted start/end indices, and none of the indices between them. Non-parenthesised values here allow hits on indices within the span. All remaining unaccounted examples in the totals were skipped due to exceeding the model max token length (512).
Model EM F1 top-1 (L) top-1 (P) top-2 (P) IoU (P) Examples Skipped
RoBERTa-L 82.29 86.27 93.70 97.47 96.22 (88.95) 92.70 (80.09) 5844 45
Flan-T5-L 83.65 90.15 95.06 98.64 94.73 (85.64) 89.94 (74.86) 5830 36
Figure A.2: Plotting the LeRF - MoRF difference in label confidence of Llama3.2-1B on IMDB classification in (a) and Wikipedia next-token prediction in (b). A higher curve is better, signifying bigger gaps in prediction confidence when removing the features with highest vs. lowest importance.
Figure A.3: Comparing faithfulness of methods using ABPC from pretrained ViT-b-16 performance on 1000 examples from the ImageNet CIFAR-10 test set. Occlusion is applied by replacing 16x16 patches with the corresponding regions from a Gaussian-blurred version of the image (kernel size = 51, $\sigma$ = 20). Red = MoRF curve, Blue = LeRF curve.
Table A.4: DynamicLRP operation coverage breakdown of 15 models spanning vision, text, and audio modalities. Unique Ops indicates the number of distinct operation types in each model’s computation graph.
Model Modality Covered Nodes Unique Ops covered
VGG16 Simonyan and Zisserman (2015) Vision 90/90 9/9
ResNet-50 He et al. (2015) Vision 339/339 10/10
ViT-b-16 Dosovitskiy et al. (2021) Vision 779/779 16/16
EfficientNetv2-medium Tan and Le (2021) Vision 1,526/1,526 12/12
Wav2Vec2-xls-r-300m Babu et al. (2021) Audio 2,021/2,022 18/19
whisper-small Radford et al. (2022) Audio 2,729/2,729 22/22
GPT-2 Radford et al. (2019) Language 885/885 19/19
RoBERTa-large Liu et al. (2019) Language 1,461/1,461 14/14
Llama3.2-1B Grattafiori et al. (2024) Language 1,787/1,787 24/24
Gemma-3-270m-it Gemma Team (2025) Language 2,482/2,482 24/24
Qwen3-0.6B Qwen Team (2025) Language 2,590/2,590 18/18
Flan-T5-large Chung et al. (2022) Language 5,713/5,713 24/24
DePlot Liu et al. (2022) Vision/Language 2,863/2,863 23/23
Mamba-130m Gu and Dao (2024) State Space 3,421/3,421 26/26
Aggregate - 31,464/31,465 -
Table A.5: Uncovered operation breakdown, where applicable.
Model Uncovered Ops Node count
Wav2Vec2-xls-r-300m WeightNormInterface 1